# ConCISE: Confidence-guided Compression in Step-by-step Efficient Reasoning

### v1

@@ 3.3.2 Early Stopping
Although Confidence Injection effectively boosts the model’s confidence during reasoning, it does not fully mitigate the pattern of Termination Delay, which is caused by the significant increase in the dynamic threshold $t_{i}$ after the model generates an answer, leading the model to continue unnecessary reflections even when its confidence is already high. To address this, we design an Early Stopping mechanism based on direct estimation of the model’s internal confidence. Specifically, we construct a lightweight confidence detector to provide a quantitative proxy for the model’s internal confidence at each step after the first answer is generated.
We introduce a probing prompt (such as "So I’m") after the current reasoning context and analyze the probability distribution of generated continuations. Based on statistical analysis, we collect a set of confidence-indicative tokens $\mathcal{W}_{+}$, which reflect affirmations of high certainty. Formally, the detected confidence score $\hat{c}_{i}$ at step $i$ is calculated as:
@@
To validate the confidence detector and examine how model confidence evolves post-FAS(First Answer Step), we calculate the average detected confidence $\hat{c}_{i}$ across key reasoning stages, including FAS and the subsequent five reflection steps (Ref1 to Ref5). As shown in Figure 3(b), the model exhibits low confidence at the FAS, but after the first reflection, confidence increases sharply and continues to rise through subsequent reflections. Notably, the final reflection step has an average detected confidence of $0.93$, suggesting that for $i>i^{\prime}$, the dynamic threshold $t_{i}$ effectively stabilizes around this high value, motivating the introduction of a manually controlled lower threshold. Therefore, we define an early stopping threshold $t_{\text{e}}=0.5$. The model is instructed to terminate generation as soon as its detected confidence $\hat{c}_{i}$ exceeds $t_{\text{e}}$. Further details regarding the selection of $t_{e}$ and the confidence detector are provided in Appendix A.2.
In addition, we also compute the average and maximum $\hat{c}_{i}$ before the FAS (Avg Pre-FAS and Max Pre-FAS), both of which remain close to zero. This indicates that our confidence detector only becomes effective after the first answer is generated, ensuring that premature termination during early reasoning stages does not occur.
@@ 3.4 Building Efficient Reasoning Chains
ConCISE integrates Confidence Injection and Early Stopping to dynamically control the reasoning generation process. Given an input question, the LRM generates the reasoning chain step-by-step. At each step, the model first examines whether the next generated step is a reflection step. If a reflection is detected, Confidence Injection is applied by appending a confidence prompt and regenerating the step with reinforced internal confidence. After generating the next reasoning step, a lightweight confidence detector estimates the detected confidence score $\hat{c}_{i}$. If $\hat{c}_{i}$ exceeds a predefined threshold $t_{e}$, the model terminates generation, assuming sufficient confidence in the answer. Otherwise, the generated step is appended to the current reasoning chain, and the process repeats.. Algorithm 1 summarizes the overall coordinated reasoning process in ConCISE.
Algorithm 1 Algorithmic workflow of ConCISE for generating concise reasoning chains.
1:Preparation:
2:    Large Reasoning Model $\pi_{\theta}$
3:    Confidence Phrase Pool $\mathcal{P}$
4:    Early Stopping Threshold $t_{e}=0.5$
5:    Prompt Template $T$
6:Input: Question $Q$
7:Initialize: Reasoning chain $S_{0}=T(Q)$
8:for each reasoning step $i=1,2,\ldots$ do
9:     Generate $s_{i}=\pi_{\theta}(S_{i-1})$
10:     if $s_{i}$ is a reflection step then
11:         Apply Confidence Injection:
12:             Sample a confidence phrase $p_{i}\in\mathcal{P}$
13:             Regenerate $s_{i}=\pi_{\theta}(S_{i-1}+p_{i})$
14:     end if
15:     Compute detected confidence $\hat{c}_{i}$
16:     if $\hat{c}_{i}>t_{e}$ then
17:         Terminate
18:     end if
19:     Accept $s_{i}$ and update $S_{i}=S_{i-1}+s_{i}$
20:end for
@@ 3.5 Training Objective and Fine-tuning
To align LRM generation with the behavior encouraged by ConCISE, we first construct a fine-tuning dataset by applying the ConCISE pipeline to a set of reasoning questions. This yields concise reasoning traces in which redundant reflections are actively suppressed and timely termination is encouraged through the use of Confidence Injection and Early Stopping.
We fine-tune the model on this dataset using two standard learning paradigms: supervised fine-tuning (SFT), which directly imitates the revised reasoning chains, and Simple Preference Optimization (SimPO), which further amplifies the model’s preference for confident, efficient reasoning. Both approaches optimize a shared objective that governs the desired generation behavior:
@@
This objective encourages the model to learn a generation policy guided by internal confidence, rather than merely imitating surface text patterns. Specifically, when the model’s confidence $c_{i}$ at a given step falls below the threshold $t_{i}$, it learns to approximate the distribution conditioned on an injected confidence phrase $p_{i}$, which helps improve internal certainty and suppress unnecessary reflections. After the first answer is generated (at step $i^{\prime}$), if the detected confidence score $\hat{c}_{i}$ exceeds the early stopping threshold $t_{e}$, the model is explicitly trained to terminate reasoning at that point.
Through this process, the model learns to regulate its generation trajectory based on confidence, strengthening certainty during reasoning and halting once sufficient confidence is reached. As a result, the generation distribution favors precise and concise reasoning chains without compromising task performance.
@@ 4 Experiments
Model Method Math-500 GSM8K AIME24 GPQA Average
[…]
DS-7B Origin 90.8 3854.2 100% 93.1 1441.9 100% 54.2 13573.5 100% 51.0 8141.8 100% 72.3 100%
OverThinkSFT 92.2 2537.7 65.8% 93.0 1002.1 69.5% 52.5 11225.3 82.7% 51.1 7638.7 93.8% 72.2 78.0%
SpiritSFT 91.0 2934.6 76.1% 91.4 1107.4 76.8% 51.7 11528.7 84.9% 50.3 7084.3 87.0% 71.1 81.2%
ConCISESFT 92.0 2244.3 58.2% 92.9 831.9 57.7% 52.1 9750.8 71.8% 50.0 5891.5 72.4% 71.8 65.0%
OverThinkSimPO 91.4 2405.1 62.4% 92.9 879.4 61.0% 50.0 9603.2 70.8% 49.9 6304.9 77.4% 71.0 67.9%
SpiritSimPO 87.2 1765.1 45.8% 90.8 687.6 47.7% 38.3 6926.0 51.0% 50.2 5832.0 71.6% 66.6 54.0%
ConCISESimPO 91.0 1945.7 50.5% 92.1 715.1 49.6% 48.3 7744.7 57.1% 48.0 4859.4 59.6% 70.0 54.2%
DS-1.5B Origin 82.2 4783.7 100% 85.4 2219.0 100% 29.2 17464.9 100% 34.5 9492.3 100% 57.8 100%
OverThinkSFT 84.0 3295.6 68.9% 85.2 1199.7 54.1% 30.0 12893.3 73.8% 34.8 8600.0 90.6% 58.5 71.8%
SpiritSFT 83.8 3856.7 80.6% 84.8 1475.9 66.5% 28.8 13684.5 78.4% 32.3 8554.7 90.1% 57.4 78.9%
ConCISESFT 83.6 2700.8 56.5% 84.9 923.3 41.6% 30.0 11359.0 65.0% 35.0 7252.6 76.4% 58.4 59.9%
OverThinkSimPO 83.6 2737.8 57.2% 84.8 1002.9 45.2% 31.26 11465.1 65.7% 34.7 8249.8 86.9% 58.6 63.8%
SpiritSimPO 82.6 2455.3 51.3% 82.6 804.2 36.2% 26.66 9946.3 57.0% 33.1 6910.3 72.8% 56.2 54.3%
ConCISESimPO 83.6 2428.6 50.8% 84.3 803.0 36.2% 30.43 8810.1 50.4% 35.4 7056.1 74.3% 58.4 52.9%
Table 1: Evaluation results of OverThink, SPIRIT, and ConCISE fine-tuned via SFT or SimPO on DeepSeek-R1-Distill 7B and 1.5B, showing accuracy(Acc.), token count(Tok.), and compression ratio(CR) across four benchmarks.
Figure 4: Comparison of StepNum and StepLen across two models and four benchmarks. Stars indicate ConCISE-based methods, which achieve stronger stepwise compression than baselines while maintaining comparable accuracy.
@@ 4.1 Settings
Models. We evaluate ConCISE on two Large Reasoning Models (LRMs): DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Qwen-1.5B Guo et al. (2025). Although both models exhibit strong mathematical reasoning abilities, their substantial redundancy in reflection steps highlights the need for reasoning chain compression.
Question Set. We find that effective reasoning-chain compression can be achieved with relatively limited data. Based on this observation, we select 2,000 questions from the MATH training setHendrycks et al. (2021) for fine-tuning, ensuring that each question yields a correct answer under greedy decoding and that the set maintains diversity in difficulty levels and response lengths. Detailed analysis is provided in the AppendixB.1.
Evaluation Benchmarks. We evaluate model performance across four benchmarks: GSM8kCobbe et al. (2021), Math-500Hendrycks et al. (2021), AIME24Li et al. (2024), and GPQA_diamondRein et al. (2024). All experiments use the same decoding configuration with temperature = 0.6 and top_p = 0.95Guo et al. (2025). The maximum length is set to 16k for GSM8k, Math-500, GPQA_diamond, and 32k for AIME24. For AIME24 and GPQA_diamond, due to their higher difficulty and smaller sizes, we perform 8 sampling rounds per instance and report the mean values across all metrics to ensure stability and reliability.
Evaluation Metrics. We adopt three primary metrics to comprehensively assess both reasoning accuracy and compression effectiveness. Acc. denotes the final answer accuracy. Tok. refers to the average response length, measured in tokens. CR (Compression Rate) is defined as the ratio of the average response length to that of the original model, with lower values indicating higher compression.
Baseline Methods. We compare our method against existing approaches that aim to remove redundant reasoning and construct efficient reasoning datasets. OverThink.Chen et al. (2024) Samples eight responses from the LRM with a relatively high temperature and selects the shortest correct one. Only the first reflection step after the answer is retained. Spirit.Cui et al. (2025) Iteratively removes steps with minimal impact on perplexity (PPL), deleting 30% of the least critical steps. Adjacent contexts are merged when needed to maintain coherence.
Training Details. We construct training datasets based on the question set using ConCISE, OverThink, and Spirit. For Spirit, compression is applied to reasoning chains generated via greedy search. During training, we adopt two strategies: SFT and SimPO. SFT directly uses the question and its corresponding compressed reasoning trace for supervised fine-tuning. For SimPO, reject samples are also required Rafailov et al. (2023). Following the original OverThink setup, we use the longest correct reasoning chain among the eight samples as the reject samples. To ensure a fair comparison, ConCISE adopts the same strategy. For Spirit, the full reasoning chain before compression serves as reject samples. All models are trained with $lr=1e-6$; SFT runs for 2 epochs and SimPO for 1 epoch.
@@ 4.2 Reults and Analysis
This section evaluates multiple compression methods across various reasoning benchmarks. Our results show that ConCISE strikes a superior balance between compression and task performance compared to baseline methods, enabling models to effectively eliminate redundant reasoning steps. Additionally, models fine-tuned with ConCISE on mathematical datasets generalize well to GPQA_diamond, showcasing its robustness.
@@
4.2.1 ConCISE achieves superior compression-performance balance
As shown in Table 1, ConCISE demonstrates strong reasoning chain compression under both SFT and SimPO settings while maintaining comparable accuracy. Under SFT, it achieves a compression rate of ~60%, and under SimPO, it further improves compression to ~50%.
@@ 4.2.1 ConCISE achieves superior compression-performance balance
In contrast, OverThink achieves competitive task performance but exhibits weaker compression under SFT. Although it achieves further compression under SimPO, its final compression rate and task performance are comparable to, or even slightly worse than, ConCISE under SFT. As for Spirit, it underperforms in both settings, achieving the weakest compression under SFT and suffering noticeable task performance degradation compared to the other two methods. While Spirit matches ConCISE’s compression rate under SimPO, its accuracy drops more severely, especially on challenging datasets like AIME24. Overall, ConCISE offers a better balance between compression and performance compared to baseline methods.
(a) Step-level analysis of reasoning chains on Math-500
(b) Comparison of training datasets of different methods
Figure 5: Extended analysis of reasoning chain characteristics. More details can be found in the Appendix B.3.
4.2.2 How does ConCISE achieve better compression-performance balance?
@@
To deeply understand how ConCISE achieves a better balance between compression and performance, we perform a stepwise analysis of generated reasoning chains on benchmarks. As shown in Figure 4, we evaluate the average number of reasoning steps in the thinking stage(StepNum) and the average token count per step in the thinking stage(StepLen) across different benchmarks. ConCISE consistently achieves the fewest reasoning steps, while StepLen remains similar to that of the original model. Further details can be found in the AppendixA.3.
Further analysis, focusing on reflection steps in the generated reasoning chains on MATH-500, reveals how each method handles reflection behavior. We use several metrics, including StepNum, Non-RefNum, RefNum, Pre-FAS RefNum, and Post-FAS RefNum. Non-RefNum represents the number of non-reflection steps, and the last three metrics measure the number of reflection steps across the entire reasoning chain, before FAS(First Answer Step), and after FAS, respectively. The results are shown in Figure 5(a).
We find that, in terms of RefNum, Spirit-SFT retains the most reflection steps, followed by OverThink, while ConCISE and Spirit-SimPO retain fewer reflection steps. For Non-RefNum, OverThink and ConCISE perform similarly, while Spirit-SimPO slightly underperforms and also causes a performance drop. This suggests that none of the compression methods can effectively reduce non-reflection steps without compromising model performance. Consequently, it indicates that existing compression methods mainly achieve compression by pruning reflection steps.
For Pre-FAS RefNum, OverThink scores the highest, followed by Spirit, as neither method effectively controls early-stage reasoning. Spirit-SFT, on the other hand, has the highest Post-FAS RefNum, indicating its limited ability to identify and prune redundant steps. In contrast, ConCISE maintains consistent control over both pre- and post-answer reflections. These differences help explain the performance gaps: OverThink’s inability to prune pre-answer reflections limits its compression potential, while Spirit’s failure to identify redundant steps leads to suboptimal pruning, removing essential content and degrading performance. ConCISE, through confidence injection and early stopping, precisely removes redundant reflection steps without harming essential reasoning process.
4.2.3 How does the training dataset affect compression?
Since all methods share identical training configurations and hyperparameters, the quality of the training dataset directly determines the final compression performance. To further investigate which features of the training dataset are highly correlated with the final compression performance, we analyzed the reasoning datasets constructed by ConCISE, OverThink, Spirit, Greedy-Search (reject samples of Spirit), and Longest-of-N (reject samples of ConCISE and OverThink), with the evaluation metrics including Tok., StepNum, StepLen, RefNum, Pre-FAS RefNum, and Post-FAS RefNum. Figure 5(b) shows the distribution of these metrics across the training datasets.
Our results indicate that surface-level metrics such as Tok., StepNum, and StepLen do not directly correlate with the compression rate. For instance, ConCISE and OverThink have nearly identical values on these metrics, but ConCISE achieves much better compression after fine-tuning.
In contrast, other metrics reveal clear distinctions: OverThink records the highest Pre-FAS RefNum, while Spirit leads in Post-FAS RefNum. ConCISE, by comparison, consistently maintains lower values on both. This aligns with the reasoning chain analysis in the previous subsection, indicating that key features from the training data are effectively transferred to the fine-tuned model. These features enable the model to compress reasoning, particularly reflection steps, in line with the dataset design principles.
4.2.4 Can reasoning compression methods generalize across tasks?
Another important question is whether compression methods can generalize beyond the training data distribution. This property is crucial for the practical use of reasoning chain compression.Qu et al. (2025) In our setup, the training dataset is constructed from the question set of the MATH training set. Accordingly, MATH-500 serves as the in-distribution evaluation, GSM8K and AIME24 represent in-domain but out-of-distribution tasks, and GPQA_diamond serves as an out-of-domain task, differing substantially in content, format, and prompt structure.
As shown in Table 1, ConCISE consistently achieves a better balance between compression and performance than baseline methods across both in-domain and out-of-domain tasks. These results indicate that fine-tuning on efficient reasoning chains constructed by ConCISE enables models to structurally suppress redundant reflection steps, a property that holds regardless of task type or prompt format, without compromising overall performance.

### v2

@@ 3.3.2 Early Stopping
(a) Reflection probability of next step after phrase injection.
(b) Confidence calculated by our detector.
Figure 4: Effectiveness and necessity of Confidence Injection and Termination Delay(details in Appendix B.1).
Although Confidence Injection effectively boosts the model’s confidence during reasoning, it does not fully mitigate Termination Delay, which is caused by the high threshold $t_{i}$ after the model reaches an answer, leading the model to continue unnecessary reflections even when its confidence is already high. To address this, we design an Early Stopping mechanism based on direct estimation of the model’s internal confidence.
Specifically, we construct a lightweight confidence detector to provide a quantitative proxy for the model’s internal confidence about the answer. We introduce a probing prompt after the current reasoning context and analyze the probability distribution of generated continuations. Based on the statistical analysis shown in Appendix A.2, we collect a set of confidence-indicative phrases $\mathcal{W}_{+}$, which reflect affirmations of high certainty. The detected confidence score $\hat{c}_{i}$ at step $i$ is calculated as:
@@
where $p(\cdot|\cdot)$ denotes the continuation probabilities determined by $\pi_{\theta}$. We then calculate the average $\hat{c}_{i}$ after the First Answer Step (FAS) is generated, including the subsequent five reflection steps (Ref1-5). As shown in Figure 4(b), the model exhibits relatively low confidence at the FAS, but increases sharply after the first reflection, and continues to rise through subsequent reflections. Notably, when the reasoning chain stops, the average confidence is $0.95$, motivating the introduction of a manually controlled lower threshold $t_{e}$. We ultimately set $t_{e}=0.5$ through experiments.
Additionally, the application of the threshold is carefully controlled to prevent premature termination: once the detected confidence $\hat{c}_{i}$ exceeds $t_{e}$, the model is first prompted to output a final answer (e.g., by appending Final Answer:). The reasoning process halts only if this answer is subsequently verified as correct; otherwise, generation continues. More details regarding the selection of $t_{e}$ and Early Stopping are provided in Appendix A.2.
@@ 3.4 Building Efficient Reasoning Chains
Preparation: LRM generation policy: $\pi_{\theta}$, Confidence Phrase Pool: $\mathcal{P}$, Early Stopping Threshold $t_{e}$, Prompt Template $T$
Input: Question $q$, Ground Truth gtgt
Initialize: Reasoning chain $S_{0}\leftarrow T(q)$
for each reasoning step $i=1,2,\ldots$ do
 $s_{i}\leftarrow\pi_{\theta}(S_{i-1})$
 if $s_{i}$ is a reflection step then
      Sample a confidence phrase $p_{i}\in\mathcal{P}$ $s_{i}\leftarrow\pi_{\theta}(S_{i-1},p_{i})$
 $S_{i}\leftarrow S_{i-1}+s_{i}$
   Compute detected confidence $\hat{c}_{i}$
 if $\hat{c}_{i}>t_{e}$ then
    $a\leftarrow\pi_{\theta}(S_{i},\text{'Final Answer:'})$
    if $\text{isequal}(a,gt)$ then
       break
if $\text{isequal}(a,gt)$ then
   Summary: $S\leftarrow S_{i}+\pi_{\theta}(S_{i},^{\prime}\text{</think>}^{\prime})$
 Output: ConCISE reasoning chain $S$
else
 $S_{i}$ and $q$ are discarded
Algorithm 1 Workflow of ConCISE.
Model Math-500 GSM8K AIME24 GPQA Average
[…]
DeepSeek-7BOrigin 90.8 3854 100% 93.1 1442 100% 54.2 13574 100% 51.0 8142 100% 72.3 100%
OverThinkSFT 92.2 2538 66% 93.0 1002 70% 52.5 11225 83% 51.1 7639 94% 72.2 78%
SpiritSFT 91.0 2935 76% 91.4 1107 77% 51.7 11529 85% 50.3 7084 87% 71.1 81%
ConCISESFT 92.0 2244 58% 92.9 832 58% 52.1 9751 72% 50.0 5892 72% 71.8 65%
OverThinkSimPO 91.4 2405 62% 92.9 879 61% 50.0 9603 71% 49.9 6305 77% 71.0 68%
SpiritSimPO 87.2 1765 46% 90.8 688 48% 38.3 6926 51% 50.2 5832 72% 66.6 54%
ConCISESimPO 91.0 1946 51% 92.1 715 50% 48.3 7745 57% 48.0 4859 60% 70.0 54%
DeepSeek-1.5BOrigin 82.2 4784 100% 85.4 2219 100% 29.2 17465 100% 34.5 9492 100% 57.8 100%
OverThinkSFT 84.0 3296 69% 85.2 1200 54% 30.0 12893 74% 34.8 8600 91% 58.5 72%
SpiritSFT 83.8 3857 81% 84.8 1476 67% 28.8 13685 78% 32.3 8555 90% 57.4 79%
ConCISESFT 83.6 2701 57% 84.9 923 42% 30.0 11359 65% 35.0 7253 76% 58.4 60%
OverThinkSimPO 83.6 2738 57% 84.8 1003 45% 31.3 11465 66% 34.7 8250 87% 58.6 64%
SpiritSimPO 82.6 2455 51% 82.6 804 36% 26.7 9946 57% 33.1 6910 73% 56.2 54%
ConCISESimPO 83.6 2429 51% 84.3 803 36% 30.4 8810 50% 35.4 7056 74% 58.4 53%
Skywork-7BOrigin 93.6 4178 100% 93.2 2111 100% 62.9 12464 100% 51.1 8374 100% 75.2 100%
OverThinkSFT 92.8 3410 82% 93.1 1561 74% 59.2 11047 89% 50.7 8128 97% 73.9 85%
SpiritSFT 93.4 3279 79% 93.1 1452 69% 58.8 11815 95% 52.0 7565 90% 74.2 83%
ConCISESFT 93.2 2740 66% 92.7 1247 59% 59.2 9871 79% 51.8 6543 78% 74.2 71%
OverThinkSimPO 93.8 2758 66% 93.4 1114 53% 56.3 9422 76% 51.5 6773 81% 73.7 69%
SpiritSimPO 92.0 2329 56% 92.3 801 38% 52.5 8055 65% 51.0 6663 80% 72.0 59%
ConCISESimPO 93.0 2207 53% 93.3 882 42% 55.9 7598 61% 51.7 5668 68% 74.0 56%
Qwen3-8BOrigin 93.4 5142 100% 95.6 2211 100% 73.3 15094 100% 60.6 7309 100% 80.7 100%
OverThinkSFT 93.2 4963 97% 95.5 2133 96% 72.5 14123 94% 60.3 7204 99% 80.4 96%
SpiritSFT 93.2 4184 81% 95.2 1675 76% 73.3 13854 92% 60.0 6784 93% 80.4 85%
ConCISESFT 93.0 4712 92% 95.6 1907 86% 72.1 14168 94% 60.4 6952 95% 80.6 92%
OverThinkSimPO 93.6 2488 48% 95.2 929 42% 69.2 9468 63% 57.1 3524 48% 78.9 50%
SpiritSimPO 93.2 2920 57% 95.3 997 45% 67.9 10441 69% 56.8 3473 48% 78.3 55%
ConCISESimPO 93.0 2271 44% 95.5 841 38% 71.6 10098 67% 57.6 3751 51% 79.4 50%
Table 1: Comparison of OverThink, SPIRIT, and ConCISE methods, fine-tuned with SFT or SimPO on four LRMs. Metrics include Accuracy (Acc.), Token Count (Tok.), and Compression Ratio (CR) over four benchmarks. Light gray rows denote original LRM performance. Best and second-best results are in bold and underlined respectively. Note that the “Average” column is computed as an unweighted mean across benchmarks, and this convention applies to all subsequent tables as well.
@@
ConCISE integrates Confidence Injection and Early Stopping to dynamically control the reasoning generation process. Given an input question, the LRM generates the reasoning chain step-by-step. At each step $s_{i}$: if it is identified as a reflection step, Confidence Injection is applied by appending a confidence prompt and regenerating $s_{i}$ with boosted internal confidence. $s_{i}$ is then added to the current chain $S_{i}$. Subsequently, a lightweight confidence detector estimates the detected confidence score $\hat{c}_{i}$. If $\hat{c}_{i}>t_{e}$, the LRM is prompted to provide a final answer $a$. Crucially, the generation process terminates early at this stage if, and only if, this answer $a$ matches the ground truth $gt$ and $\hat{c}_{i}>t_{e}$, otherwise reasoning continues to the next step. After the loop concludes, should a correct final answer $a$ have been verified, a concluding summary (e.g., prompted by </think>) is generated and appended to the accumulated steps $S_{i}$ to form the complete ConCISE reasoning chain $S$, otherwise, if a wrong answer was finally achieved, the partial chain and the original question are discarded. Algorithm 1 summarizes this coordinated reasoning process in ConCISE.
@@ 3.5 Training Objective and Fine-tuning
To align LRM generation with the behavior encouraged by ConCISE, we first construct a fine-tuning dataset by applying the ConCISE pipeline to a set of questions. This yields concise reasoning chains in which redundant reflections are actively suppressed. We then fine-tune the LRM on this dataset using two standard learning paradigms: Supervised Fine-Tuning (SFT) and Simple Preference Optimization (SimPO). Both approaches optimize a shared objective that governs the desired generation behavior:
@@
Specifically, when the model’s confidence $c_{i}$ at a given step falls below the threshold $t_{i}$, it learns to approximate the distribution conditioned on an injected confidence phrase $p_{i}$, which helps improve internal certainty and suppress unnecessary reflections. After the first answer is generated, if the detected confidence score $\hat{c}_{i}$ exceeds the early stopping threshold $t_{e}$, the model is explicitly trained to terminate reasoning at that point. Through fine-tuning, the model could learn to regulate its generation trajectory based on confidence, strengthening certainty during reasoning and halting once sufficient confidence is reached.
(a) Average Acc, StepNum, and StepLen.
(b) Analysis of Reflection steps.
(c) Metrics of training datasets.
Figure 5: Further analysis of reasoning chains and training datasets on DeepSeek-R1-Distill-Qwen-7B.
@@ 4.1 Settings
Models and Datasets.
We evaluate ConCISE on four LRMs: DeepSeek-R1-Distill-Qwen-7B & 1.5BGuo et al. (2025), Skywork-OR1-7B-PreviewHe et al. (2025), and Qwen3-8BTeam (2025). As for the training dataset, we select 2,000 questions from the MATH training setHendrycks et al. (2021), ensuring that each question yields a correct answer under greedy decoding and maintaining diversity in difficulty and response lengths. Importantly, for each model, we generate its own fine-tuning data using the ConCISE pipeline, rather than relying on external models, which ensures consistency between the training data and the reasoning style of the target model.
Evaluation.
We evaluate model performance across four benchmarks: GSM8KCobbe et al. (2021), Math-500Hendrycks et al. (2021), AIME24, and GPQA_diamondRein et al. (2024). All evaluations use the same decoding configuration with temperature = 0.6 and top_p = 0.95Guo et al. (2025). The maximum length is set to 16k for GSM8k, Math-500, GPQA_diamond, and 32k for AIME24. For AIME24 and GPQA_diamond, due to their higher difficulty and smaller sizes, we sample 8 times and report the mean values. As for metrics, we adopt three primary metrics to assess both reasoning accuracy and compression effectiveness comprehensively. Acc. denotes the accuracy of the final answer. Tok. refers to the average response length, measured in tokens. CR (Compression Rate) is defined as the ratio of the average response length to that of the original model, with lower values indicating better compression.
Baselines.
We compare our method against existing approaches that aim to remove redundant reasoning and construct efficient reasoning datasets, with further comparisons after fine-tuning under SFT and SimPO scenarios. OverThink Chen et al. (2024) samples eight responses from the LRM with a relatively high temperature and selects the shortest correct one. Only the first reflection step after the answer is retained. Spirit Cui et al. (2025) iteratively removes steps with minimal impact on perplexity (PPL), deleting 30% of the least critical steps. Adjacent contexts are merged to maintain coherence. Appendix B.2 shows more implementation and training details.
@@ 4.2 Reults and Analysis
This section evaluates different compression methods across various reasoning benchmarks. Our results show that ConCISE strikes a superior balance between compression and task performance compared to baseline methods, enabling models to effectively eliminate redundant reasoning steps. Additionally, models fine-tuned with ConCISE on mathematical datasets generalize well to GPQA_diamond, showcasing its robustness.
@@
ConCISE achieves superior balance between compression and task performance
As shown in Table 1, ConCISE consistently demonstrates strong compression ability across all four LRMs and under both SFT and SimPO settings, while maintaining model performance. Particularly under SimPO settings, ConCISE achieves a compression rate of ~50% with minimal impact on the original model’s performance, achieving excellent compression results on the challenging task AIME24 as well as the out-of-domain task GPQA.
@@
In contrast, OverThink achieves competitive task performance but exhibits weaker compression under both SFT and SimPO settings. Regarding Spirit, despite attaining a compression rate comparable to that of ConCISE under SimPO settings, its task performance suffered a notable degradation, especially on challenging tasks like AIME24. Overall, compared to baseline methods, ConCISE offers a better balance between compression and performance, demonstrating excellent in-domain and out-of-domain generalization and robustness.
How does ConCISE achieve a better compression & performance balance?
@@
To better understand how ConCISE achieves an improved balance between compression and performance, we particularly analyze the structure of reasoning chains generated on various benchmarks of DeepSeek-R1-Distill-Qwen-7B. As shown in Figure 5(a), we evaluate the average number of steps (StepNum) and the average token count per step (StepLen). ConCISE generates the fewest reasoning steps, while its StepLen remains comparable to that of the original model.
More detailed analysis focuses on reflection steps within reasoning chains on MATH-500, revealing distinct compression strategies among the methods. As shown in Figure 5(b), we find that pruning reflection steps is key to compression, and preserving non-reflection steps is paramount for model performance. ConCISE uniquely achieves this critical balance. Specifically, Spirit-SFT retains the most reflection steps (RefNum), followed by OverThink, while ConCISE retains the fewest. Regarding non-reflection steps (Non-RefNum), OverThink and ConCISE exhibit similar counts. Spirit shows slightly fewer non-reflection steps under SimPO, albeit at the cost of a performance drop. This suggests that current compression techniques primarily achieve compression by pruning reflection steps, as none of the evaluated methods can effectively reduce non-reflection steps without impairing model performance.
Further dissecting the reflection steps, OverThink and Spirit have the most reflection steps before FAS(Pre-FAS RefNum) due to ineffective early-stage reasoning control. In contrast, Spirit-SFT has most reflection steps after FAS(Post-FAS RefNum), indicating its inability to prune redundant post-answer reflections. These distinct strategies in reflection management contribute to performance disparities: OverThink’s limited pre-answer pruning curtails its compression potential, while Spirit’s difficulty in accurately identifying redundant steps often degrades performance by removing essential content. ConCISE, however, achieves a better compression-performance balance by precisely eliminating redundant reflection steps throughout the entire reasoning chain without compromising the essential reasoning process. We present more details about the above analysis in Appendix B
How does the training dataset affect compression?
Since all methods operate under identical training configurations, variations in final compression performance are fundamentally rooted in the characteristics of their respective training datasets. To this end, we analyze training datasets constructed by ConCISE, OverThink, Spirit, Greedy-Search (derived from Spirit’s rejected samples), and Longest-of-N (derived from rejected samples of ConCISE and OverThink).
As shown in Figure 5(c), surface-level metrics such as response tokens (Tok.), step count (StepNum), and average step length (StepLen) show no strong direct correlation with the final compression rate. For instance, while the training datasets of ConCISE and OverThink yield nearly identical values for these metrics, ConCISE demonstrates significantly better compression after fine-tuning under both SFT and SimPO. Conversely, Metrics related to reflection offer more insightful distinctions among the methods. For instance, OverThink has the highest values for Pre-FAS RefNum, while Spirit shows the highest Post-FAS RefNum. ConCISE, in contrast, consistently maintains lower figures for both these metrics, underscoring its more effective control over the model’s reflective tendencies. Echoing the analysis from the previous subsection, these observed characteristics within the training datasets directly mirror the fine-tuned model’s capacity for reflection management. This provides strong evidence that the model effectively learns features about reflection from the training data during the fine-tuning process. Thus, it becomes clear how ConCISE uses its unique training data design to achieve significant compression by identifying and cutting down on unnecessary reflection steps.
4.3 Ablation Study